OpenStudy (anonymous):

How would I simplify -2sin^2x+sinx=-1 to the point where I could solve.

OpenStudy (anonymous):

\[-2\sin^2x+sinx=-1\]

OpenStudy (anonymous):

try factoring out sinx, \[\sin x (1-2\sin x)=-1\] then use trig identity sinx(cos 2x)=-1 then i don't know what's next, sorry..

OpenStudy (anonymous):

ty